Lily’s Loving Laundry is thinking about installing a new steam press at its facility so it can add in-house pressing services to its product offerings. The new press steamer will cost $14,400. It will last for five years, be depreciated on a straight-line basis to a salvage value of $4000 but will be resold for $1,400 after five years (assume no taxes apply).

Lily estimates its demand to press shirts as: 9000, 9000, 9000, 9200, 9200 in years one through five respectively. A new steam press operator will need to be hired at $13.00 per hour and will work 40 hours per month for the year. That employee will get a raise of $0.25 per year every year starting in the second year. Shirts currently cost customers $3 for dry cleaning; and an additional charge of $1.35 will be added for pressing. Lily will also need to add a $1,500 annual insurance policy to cover any steam-related accidents.

Assume Lily pays 17% in tax on its taxable income and is looking at a five-year return period over which she would expect to earn a 10% rate of return.

What is the amount of Net Income in Year 1 (rounded to the nearest dollar)?

A) $2,100
B) $2,205
C) $2,000
D) None of the above
Answer #1

Solution:-

Depreciation per year (year 1)= (Initial cost-salvage value)/useful life= (14,400-4,000)/5= $2,080.

Calculation of annual steam press operator cost (year 1)= (Wage per hour*no. of hours per month*12)= (13*40*12)= $6,240

Revenue for first year= No. of shirts pressed*pressing charge per shirt= 9,000*1.35= $12,150

(Note: Since, the question is about pressing segment and no details have been given for laundry segment, the revenue for $3 per shirt for laundry should not be considered)

Pre-tax income for first year= Revenue-annual operator cost-depreciation-insurance cost= 12,150-6,240-2,080-1,500= $2,330

Net income year 1 (profit after taxes)= Pre-tax profits*(1-tax rate)= 2,330*(1-17%)= $1,934

Since, the net income as calculated above doesn't fall in options A, B and C, therefore, the correct option is option D, i.e. None of the above.